Suppose Cameron eams $80,000 per year but must pay income taxes according to the tax schedule given in the table to the right TaxRate (%) to $5,001 to $31,000 $31,001 to $67,000 $67,001 to $143,000 $143,001 and up According to the chart, the income tax Cameron faces is a tax
Answer #1

Cameron’s yearly income is $ 80,000. Let us now calculate the total income tax Cameron would face on this income.

                     For the first bracket, $0 to $5,000 = Cameron need to pay 10% tax

                                                                                i.e. he will pay 10% of 5000 = $500

                    For the Second bracket, $5,001 to $31,000 = Cameron need to pay 15% tax

                                                                                i.e. he will pay 15% of 26,000 (Here 5,001 to 31,000 is 26,000)

                                                                                              > (15/100) x 26,000 = $ 3,900

                                                                               

                 For the Third bracket, $31,001 to $67,000 = Cameron need to pay 25% tax

                                                                                i.e. he will pay 25% of 36,000 (Here 31,001 to 67,000 is 36,000)

                                                                                              > (25/100) x 36,000 = $ 9,000

                  For the Fourth bracket, $67,001 to $80,000 = Cameron need to pay 30% tax

                                                                                i.e. he will pay 30% of 13,000 (Here 67,001 to 80,000 is 13,000)

                                                                                              > (30/100) x 13,000 = $ 3,900

Cameron’s income being $80,000, he partially falls under the fourth bracket of income tax and does not fall under the fifth bracket of income tax.

Therefore, the total tax the Cameron faces is = $500 + $3,900 + $9,000 + $3,900

                                                                                 = $ 17,300